My least favorite because I do not understand these:(
What is the focal point of 1/3(y-2)=2(^-x+5)

Answer by stanbon(75887)

My least favorite because I do not understand these:(
What is the focal point of 1/3(y-2)=2^(-x+5)
y = 3(2^(-x+5))+2
-----------
The focal point of y = 2^x is (0,1)
Your function involves the following moves of the focal:
-x+5=0 when x = 5; this moves the focal 5 to the right: (5,1)
The 3 stretches the graph so focal moves to (5,3)
The +2 raises the graph so focal ends up at (5,5)
-------------
Cheers,
Stan H.